You are planning on buying $100,000 face value of Australian Commonwealth Government Bonds. The bonds mature on 15 February 2022 and have a coupon rate of 4.75%. If your purchase will settle on 27 April 2012, and the quoted yield for the bond is 4.37%, what is the cash price of the bonds to the nearest dollar?
Answer #1

Interest amount=Face value*Coupon rate=100000*4.75%=4750
Quoted yield for the bond=4.37%
=>(Bond Price)*(4.37%)=4750
=>Bond Price=4750/4.37%=108695.6522 or $108695 (Rounded to the nearest dollar)